OpenStudy (anonymous):

Evaluate the integral of the quotient of cosine x and cosine cubed of x, dx.

OpenStudy (anonymous):

https://gyazo.com/adb16c3d8843d66bbf542fe57c60755a

OpenStudy (anonymous):

@jim_thompson5910 Chose none of these since I did \[\int\limits_{}^{}\frac{ \cos(x) }{ \cos^3(x) }\] equals \[\int\limits_{}^{} \frac{ 1 }{ \cos^2(x)}*\frac{ \cos(x) }{ \cos(x) }\]

OpenStudy (anonymous):

Which would be the integral of sec^2 (x)

jimthompson5910 (jim_thompson5910):

and what is the integral of sec^2 ?

OpenStudy (anonymous):

Wait just noticed that equals tan(x)

jimthompson5910 (jim_thompson5910):

yep since y = tan(x) dy/dx = sec^2(x)

OpenStudy (anonymous):

Got a hundred :)